A boy is 5 years older than his sister and 1/6 as old as his father. The sum of there ages is 51. How old is father, son and daughter.
Answer by opeodu(24)

Let the son's, sister's and father's age be x,y and z respectively. X=5+y---(1) x=(1/6)z 6x=z---(2) x+y+z=51---(3) from (1) y=x-5---(4) substitute (4) n (2) into (3) x+x-5+6x=51 8x=56 x=7 y=7-5=2 z=6*7=42